Automotive trades: What length on the circumference would a 20 degree angle span on a 16-in. fly wheel?
Found 2 solutions by stanbon, chitra:
Answer by stanbon(57400)

What length on the circumference would a 20 degree angle span on a 16-in. fly wheel?
--------
Circumference = (pi)d = 16(pi) inches
----
A 20 degree span will be (20/360)*16(pi) = 2.7925... inches
============
Cheers,
Stan H.

Answer by chitra(359)

The length of the circumference s = r* theta

where s is the lenght of the circumference

r is the total span of the curve

theta is the angle made

so substituing in the above formula we get:

S = 16 * 20

s = 320
thus the solution